OpenStudy (jennychan12):

How to solve integral of ((x^2+5)/(x^3-x^2+x+3))dx?

OpenStudy (jennychan12):

\[\int\limits \frac{ x^2+5 }{ x^3-x^2+x+3 }dx\] sorry, i made a typo in the original question.

OpenStudy (jennychan12):

more specifically, how would you split that into partial fractions?

OpenStudy (jennychan12):

i think i'm just a little confused on how to split up the denominator.

OpenStudy (anonymous):

i think by some miracle the denominator factors

OpenStudy (anonymous):

try \[(x^2-2x+3)(x+1)\]

OpenStudy (jennychan12):

how did you get that from the denominator? cuz i got (x^2-x+3)(x-1)

OpenStudy (jennychan12):

whoops, i meant (x^2+x+3)(x-1)

OpenStudy (anonymous):

easy, i cheated http://www.wolframalpha.com/input/?i=+ \frac{+x^2%2B5+}{+x^3-x^2%2Bx%2B3+}

OpenStudy (anonymous):

it can't be \(x-1\) because1 is not a zero of the denominator, \(-1\) is

OpenStudy (anonymous):

so it must be \(x+1\) as one factor you can get the other factor by division

OpenStudy (jennychan12):

cuz i did x^2(x-1)+x+3 which becomes (x^2+x+3)(x-1)

OpenStudy (anonymous):

can't be

OpenStudy (anonymous):

at the risk of repeating myself, \(-1\) is a zero of the denominator so it MUST factor as \((x+1)(something)\)

OpenStudy (jennychan12):

ASDFGHJKL./;" i think i just figured it out. i think.

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

cheat and use wolfram the problem was cooked so that the numerators are integers when you get partial fractions

OpenStudy (jennychan12):

cooked?

OpenStudy (jennychan12):

i got the answer anyways.

OpenStudy (anonymous):

@jennychan12 Was\[\log (x+1)+\sqrt{2} \tan ^{-1}\left(\frac{x-1}{\sqrt{2}}\right) \]your answer?

OpenStudy (jennychan12):

i think so; i don't really remember it cuz i turned in the paper. thanks tho! @robtobey